How much 12% indicator solution needs to be mixed with 16% indicator solution to het 10 gal of 14% solution?
Answer by TimothyLamb(4379)

w + s = 10
s = 10 - w
12w + 16s = 14(10)
12w + 16s = 140
---
12w + 16(10 - w) = 140
12w + 160 - 16w = 140
4w = 20
---
w = 5 gal
s = 5 gal
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php